Dear Collaborators, Here is a short-term schedule for the machine. - Physics run started one week ago (1/11). Delivered luminosity so far ~1.7 nb^-1 for STAR and PHENIX and ~0.6 nb^-1 for BRAHMS and PHOBOS. This is already ~20% of the minimum luminosity goal. - Our typical ZDC rates at the beginning of store ~6-7K Hz. "The store of the week" was on Sunday afternoon: 15KHz at the beginning of the store. - Store length is fixed at 5 hours for now. This is not so unreasonable since the lifetime of the Cu beam is longer than Au. - Beam Experiment tonight (1/18): 10 pm-6am followed by a maintenance day. 18 hours (6am 1/19 -12am 1/20) for STAR 10 hours (6am 1/19 - 4 pm 1/19) for the others. - Beam/Luminosity development during the day time for 9 am - 5 pm at least until next week. JH _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Tue Jan 18 14:59:50 2005
This archive was generated by hypermail 2.1.8 : Tue Jan 18 2005 - 15:00:03 EST